OpenStudy (anonymous):

Factor this expression. x^2 + 9x + 8

OpenStudy (anonymous):

which equation represents the line that passes thou the point (1,5) and had a slope of (-2)

OpenStudy (anonymous):

1- y= 2x+72- y= -2x+113- y=2x-94-y=2x+3

OpenStudy (anonymous):

I think it is y=2x+72?